Q Corporation and R Inc. are two companies with very similar characteristics. The only difference between the two companies is that Q Corp. is an unlevered firm, and R Inc. is a levered firm with debt of $5 million and cost of debt of 10%. Both companies have earnings before interest and taxes (EBIT) of $2 million and a marginal corporate tax rate of 40%. Q Corp. has a cost of capital of 15%.

a. What is Q’s firm value?

b. What is R’s firm value?

c. What is R’s equity value?

d. What is Q’s cost of equity capital?

e. What is R’s cost of equity capital?

f. What is Q’s WACC?

g. What is R’s WACC?

h. Compare the WACC of the two companies. What do you conclude?

i. What principle have you proven in this case?

j. Both companies are now evaluating a project that requires an initial investment of $1.15 million and that will yield cash inflows of $500,000 per year for the next three years. Assume that this project has the same risk level as the individual firm’s assets. Should Q invest in this project? Should R invest in this project?

k. Based on your results for part (j), discuss the effects of leverage and its tax shields effects on the future value of the two firms.
Answer #1

As per rules I am answering the first 4 subparts of the question

1: Value of Q = EBIT*(1-Tax)/ Required rate

= 2000000*(1-0.4)/ 15%

=8,000,000

2: Value of R= Value of Q+ Tax*Debt

= 8,000,000+ 0.4*5,000,000

=$ 10,000,000

3: Value of R’s equity = Value of R- Debt

= $10,000,000- 5,000,000

=$5,000,000

4: Q’s cost of equity capital = 15%

(This is given In the problem as cost of capital which is same as cost of equity capital since there is no debt in Q)